OpenStudy (perii224):

PLEASE HELP WILL MEDAL Evaluate a^a when a equals -2. -4 4 1/4 -1/4 A $20,000 business computer depreciates at a rate of 15% per year. Which of the following equations would model the value of the computer? y=20000(0.15)^x y=20000(0.85)^x y=20000(1.15)^x y=20000+15x

OpenStudy (perii224):

@Mehek14 Can you help?

OpenStudy (dumbcow):

when exponent is neg, take reciprocal \[(-2)^{-2} = \frac{1}{(-2)^2} = \frac{1}{4}\]

OpenStudy (dumbcow):

for an exponential model of form \[y = A(1+r)^x\] where r is rate of increase or decrease here it is decreasing by 15% so r = -.15

OpenStudy (perii224):

@dumbcow is this for the second question or first?

OpenStudy (perii224):

@welshfella can you help?

OpenStudy (dumbcow):

i answered both

OpenStudy (perii224):

Ohh okay, so the first answer is 1/4?

OpenStudy (perii224):

And then the seoncd would be..? Sorry, got confused

OpenStudy (perii224):

@dumbcow

OpenStudy (perii224):

y=20000(1.15)^x ?

OpenStudy (dumbcow):

the growth = 1+r r is rate (positive for increasing, neg for decreasing) so if function is decreasing then 1+r < 1 r = -0.15 ----> 1+r = 1-.15 = .85

OpenStudy (perii224):

so it would be y=20000(0.85)^x ?

OpenStudy (perii224):

@dumbcow

OpenStudy (perii224):

???

OpenStudy (perii224):

They were correct, thank you very much (: